OpenStudy (crashonce):

A man can go from A to B by road X or Y. The chance of a traffic jam on X is 1/5 and on Y is 1/8. If he is twice as likely to go by X as Y, chance he will hit a traffic jam is>?

OpenStudy (crashonce):

ye

OpenStudy (crashonce):

can u help with this one

OpenStudy (anonymous):

Basically, add the probability of each outcome that would lead to a traffic jam. One way is that he goes on X and gets into a traffic jam

OpenStudy (anonymous):

The other way is he goes on Y and gets into a traffic jam.

OpenStudy (anonymous):

The probability of a road having a traffic jam is independent of whether he takes it or not.

OpenStudy (anonymous):

@CrashOnce Does this help at all?

OpenStudy (crashonce):

im not sure what u mean

OpenStudy (anonymous):

What is the probability he takes road X and that he gets into a traffic jam on road X? Can you figure that out?

OpenStudy (anonymous):

Using: \[ \Pr(A\cap B) = \Pr(A)\cdot \Pr(B) \]since they are independent events.

OpenStudy (anonymous):

1. What is the probability he will take road X?

OpenStudy (crashonce):

sorry i dont get it can u explain simnpler

OpenStudy (anonymous):

"he is twice as likely to go by X as Y"

OpenStudy (anonymous):

He is only going to take one road or the other...

OpenStudy (anonymous):

If \(y\) represent the probability he takes road Y, then \(2y\) is the probability he takes road X. He must take one road or the other, so the probabilities must add up to one. This means:\[ y+2y = 1\implies 3y=1\implies y=\frac 13 \]So the probability he takes Y is \(1/3\) and t he probability he takes X is \(2/3\).

OpenStudy (anonymous):

But you aren't going to even try are you?

OpenStudy (anonymous):

Do you expect me to just give you an answer? I mean what do you know?

OpenStudy (perl):

The probability of picking road X is 2 times the probability of picking road Y Let P(Y) = k then P(X) = 2*P(Y) by substitution this gives us P(X) = 2*k since X and Y are complementary events (either you can go on road X or road Y, and must do one or the other , and can't do both) P(X) + P(Y) = 1 2k + k = 1 3k = 1 k = 1/3 therefore P(X) = 2/3 P(Y) = 1/3 We are not done though...

OpenStudy (perl):

can you solve it now?

OpenStudy (anonymous):

I think 8/13. Am I right?

OpenStudy (crashonce):

@wio i had to go so i couldnt answer you, please try not to be so rude